OpenStudy (yesgeorgie8):

if the arc length is 15 feet find theta to the nearest hundreth of a degree

OpenStudy (phi):

do they give you the radius ?

OpenStudy (yesgeorgie8):

yes

OpenStudy (anonymous):

\[\theta=\frac{ l }{ r }=\frac{ 15 }{ 6 }~radians=\frac{ 5 }{ 2 }~radians\]=?degrees

OpenStudy (anonymous):

do you know how to change in degrees?

likeabossssssss (likeabossssssss):

\(\Huge\color{#EB00FF}{\text{WELCOME}}\) \(\Huge\color{blue}{\text{TO}}\) \(\Huge\color{green}{\text{OPEN}}\)\(\Huge\color{purple }{\text{ STUDY!!!!!!!!!!!}}\) \(\Huge\heartsuit\)